OpenStudy (anonymous):

slope intercept form 3x+2y=6 divide 2y by 2 and 6 by 2 3x+y=3 3x-3=-y -y=3x/1 -3 is it right?

OpenStudy (aum):

If you divide by a constant, you should divide every term in the equation. If you divide just some and not others you will be altering the equation.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

y=3/2x - 3 is your answer

OpenStudy (aum):

3x+2y=6 We want to isolate y. Subtract 3x from both sides: 2y = 6 - 3x divide both sides by 2: y = 3 - 3/2x or y = -3/2x + 3

OpenStudy (anonymous):

Slope Intercept Form = y=mx + b So you are trying to simplify y/ get it by itself. First, subtract 3x from both sides. You now have -2y=6-3x. Now, divide each side by -2. You now have y = -3 + 3/2x. But this is not in slope-intercept form yet. You have to swap the 3/2x and the -3

OpenStudy (anonymous):

oh thanks i understand

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

they need to add 2 best answers :d